OpenStudy (anonymous):

One morning Martin drove 200 miles in 2 hours. After lunch, he covered 350 miles more in 3 hours. What was his average rate of spped in miles per hour, for the 5 hours driving?

OpenStudy (anonymous):

200+350=550 3+2=5 550/5=110 110 miles per hour may be

OpenStudy (anonymous):

110 mph